Let us denote Sn = an + bn + cn for arbitrary numbers a, b, c. It is known that S1 = 8, 5, S2 = 74, 25, S3 = 639, 625 for some values of a, b, c. What is the largest possible value of S2811 — S810S812


1
Expert's answer
2020-12-02T18:06:19-0500

"S_n = a^n + b^n +c^n"

"S_{811}^2-S_{810}\cdot S_{812}=(a^{811}+b^{811}+c^{811})^2-(a^{810}+b^{810}+c^{810})(a^{812}+b^{812}+c^{812})=a^{1622}+b^{1622}+c^{1622}+2a^{811}b^{811}+2a^{811}c^{811}+2b^{811}c^{811}-a^{1622}-b^{1622}-c^{1622}-a^{810}(b^{812}+c^{812})-b^{810}(a^{812}+c^{812})-c^{810}(a^{812}+b^{812})=2a^{811}b^{811}+2a^{811}c^{811}+2b^{811}c^{811}-a^{810}(b^{812}+c^{812})-b^{810}(a^{812}+c^{812})-c^{810}(a^{812}+b^{812})"

"S_1=8.5,\ S_2=74.25, \ S_3=639.625:"

"\begin{cases} a+b+c=8.5\\a^2+b^2+c^2=74.25\\a^3+b^3+c^3=639.625
\end{cases}"

Solving the system of equations:

"a^2+b^2+c^2=(a+b+c)^2-2(ab+bc+ac)=8.5^2-2(ab+bc+ac)"

and

"a^3+b^3+c^3=(a+b+c)(a^2+b^2+c^2-(ab+bc+ac))+3abc="

"=8.5(8.5^2-2(ab+bc+ac)-(ab+bc+ac))+3abc="

"=8.5(8.5^2-3(ab+bc+ac))+3abc"

Then we have

"\begin{cases} a+b+c=8.5\\8.5^2-2(ab+bc+ac)=74.25\\8.5(8.5^2-3(ab+bc+ac))+3abc=639.625
\end{cases} \implies"

"ab+bc+ac=-1"

"3abc=639.625-8.5(8.5^2+3)=0"

Since "abc=0" lets assume "a=0"

Then we have

"\begin{cases}b+c=8.5\\b^2+c^2=74.25\end{cases}\implies \begin{cases}b=8.5-c\\(8.5-c)^2+c^2=74.25\end{cases}"

"c^2-8.5c-1=0\implies c=\frac{17\pm\sqrt{305}}{4}"

Then "b=8.5-c=\frac{17}{2}-\frac{17\pm\sqrt{305}}{4}=\frac{17\mp\sqrt{305}}{4}"

Since the system is symmetric there is one solution:

"a=0, \ b = \frac{17-\sqrt{305}}{4}, \ c=\frac{17+\sqrt{305}}{4}"

Thus

"S_{811}^2-S_{810}\cdot S_{812}=2(\frac{17-\sqrt{305}}{4})^{811}(\frac{17+\sqrt{305}}{4})^{811}-(\frac{17-\sqrt{305}}{4})^{810}(\frac{17+\sqrt{305}}{4})^{812}-(\frac{17+\sqrt{305}}{4})^{810}(\frac{17-\sqrt{305}}{4})^{812}=-(\frac{17-\sqrt{305}}{4})^{810}(\frac{17+\sqrt{305}}{4})^{810}((\frac{17-\sqrt{305}}{4})^2+2\frac{17-\sqrt{305}}{4}\frac{17+\sqrt{305}}{4}+(\frac{17+\sqrt{305}}{4})^2)=-(\frac{289-305}{16})^{811}\cdot\frac{289}{4}=-(-1)^{811}\cdot\frac{289}{4}=72.25"


Answer: 72.25